# StableMotion: Training Motion Cleanup Models with Unpaired Corrupted Data

### v1

@@ 9. Benchmark Experiments: BrokenAMASS
Qualitative Results. As shown in Figure 9, our model effectively fixes motion artifacts such as jittering and unnatural poses across a diverse range of motions. Furthermore, as illustrated in Figure 8, when the initial frames of a jumping motion are highly corrupted, our model is able to restore the corrupted frames with a natural transition between the run and jumping phases of the motion. Our test-time techniques, including adaptive cleanup and quality-aware ensembling, can further improve cleanup performance, particularly in preserving the original motion content, as illustrated in Figure 10. Qualitative comparisons with baseline methods are available in the supplementary video.
@@ 10. Discussion and Future Work
In this work, we present StableMotion, a novel framework for training motion cleanup models directly on unpaired, corrupted motion datasets. StableMotion employs a diffusion-based generate-discriminate model that simultaneously generates motions of varying quality and predicts their quality through the proposed quality indicator variables. Using a simple masked inpainting strategy, the model effectively cleans a wide range of motion artifacts without relying on training with clean motion data, which is often costly and time-consuming to acquire. Quantitative and qualitative evaluations on SoccerMocap demonstrate StableMotion’s effectiveness in real-world production datasets.
Our method demonstrates that a clean motion dataset is not necessarily a prerequisite for developing effective motion cleanup models, opening potential opportunities of training models on large-scale, raw motion data commonly found in animation and related fields. However, the inpainting procedure utilized in our framework can lead to drastic changes of the content of a motion clip in some scenarios. For applications with higher motion precision requirements, such as dexterous manipulation, test-time classifier guidance may be incorporated to better preserve the content of the original motions. Another notable limitation of our framework is its reliance on motion quality labels, which can still involve a considerable amount of manual effort to produce. However, we show in Appendix D that effective motion cleanup models can be trained when only a small portion of the dataset is annotated with quality labels. Developing automated techniques that can accurately identify motion artifacts can further improve the effectiveness of our framework and expand its application across more diverse domains.
@@ Appendix A Evaluation Metrics Details
Figure 11. Our StableMotion model is implemented using a diffusion transformer (Peebles and Xie, 2023). The key modifications we introduce is the incorporation of rotary positional embeddings (RoPE) instead of absolute positional embeddings, which are applied to the hidden vectors $Q$ and $K$ in the multi-head self-attention module. The operation “@@@@” denotes application of a rotation to the vectors. The final key/query used for scaled dot-product attention is constructed by concatenating the original $Q/K$ with their rotated counterparts, $Q_{\text{rot}}/K_{\text{rot}}$.
Recall at Human-Annotated Artifacts (R@HAA). As described in Section 8, challenging artifacts in SoccerMocap, such as body penetration and subtle pose errors that are difficult for automatic algorithms to detect, are manually annotated by artists. We calculate our model’s artifact detection recall at these human-annotated artifacts (HAA). R@HAA is calculated following the recall definition:
[…]

### v2

@@ 9. Benchmark Experiments: BrokenAMASS
As shown in Figure 9, our model effectively fixes motion artifacts such as jittering and unnatural poses across a diverse range of motions. Furthermore, as illustrated in Figure 8, when the initial frames of a jumping motion are highly corrupted, our model is able to restore the corrupted frames with a natural transition between the running and jumping phases of the motion. Our test-time techniques, including adaptive cleanup and quality-aware ensembling, can further improve cleanup performance, particularly in preserving the original motion content, as illustrated in Figure 10. Qualitative comparisons with baseline methods are available in the supplementary video.
Figure 11. StableMotion is able to fix highly jittery motions from PopDanceSet.
10. Real-world Experiments: PopDanceSet
To further demonstrate the application of our domain-specific motion cleanup method to in-the-wild datasets, we apply StableMotion to PopDanceSet (Luo et al., 2024), a dance motion dataset collected from web videos using pose estimation algorithms. It contains over 3 hours of diverse dance motions from more than 100 subjects, recorded at 30 fps and represented in the SMPL format (Loper et al., 2015). However, due to the inherent challenges of 3D pose estimation from monocular video—video-based mocap datasets often exhibits low motion quality and contains numerous jittery artifacts, as reported in Table 3. We adopt similar evaluation metrics as the BrokenAMASS experiments. However, since clean ground-truth motions are unavailable, metrics such as Accel and GMPJPE are computed with respect to the raw input motions. To further assess the severity of different artifacts in real-world mocap data, we include additional metrics, such as Foot Skating Rate (FS Rate) and Foot Penetration Distance (FP Dist).
We compare our method against both a diffusion-based and a non-diffusion-based approach, denoted as CondMDI-B and ConvAE-B, following the implementations from the previous experiments. All models are trained on the PopDanceSet training split and evaluated on the test split. The quality labels for training StableMotion are annotated using automatic heuristics. During testing, since not all baselines are able to detect corrupted frame, we use the quality labels predicted by StableMotion for all methods to ensure a fair comparison. As shown in Table 3, despite being trained on a highly corrupted dataset, our model is able to effectively artifacts while preserving the original motion content. An example of cleanup motion jittering using StableMotion is shown in Figure 11. The non-diffusion-based method ConvAE-B is also able to effectively reduce jitter artifacts, which may be due to deterministic models being more robust to jittery artifacts that resemble Gaussian noise (Lehtinen et al., 2018). However, these deterministic models tend to struggle for more challenging dataset, such as the highly dynamic SoccerMocap dataset or BrokenAMASS, where the prevailing artifacts may not resemble Gaussian noise.
Table 3. Quantitative comparison on PopDanceSet. Results are averaged over three random seeds and reported with standard deviations. The best performing model is highlighted.
Method FS Dist $\downarrow$ FS Rate $\downarrow$ FP Dist $\downarrow$ Jitter $\downarrow$ Accel $\downarrow$ GMPJPE $\downarrow$
Input 6.41 17.62 % 12.35 27.97 0.00 0.00
ConvAE-B ${4.78}^{\pm{0.11}}$ ${12.47}^{\pm{1.98}}$ % ${8.30}^{\pm{1.04}}$ ${15.87}^{\pm{0.13}}$ ${8.51}^{\pm{0.01}}$ ${3.71}^{\pm{0.05}}$
CondMDI-B ${5.87}^{\pm{0.05}}$ ${15.28}^{\pm{0.42}}$ % ${8.56}^{\pm{0.17}}$ ${22.59}^{\pm{0.15}}$ ${9.27}^{\pm{0.03}}$ ${3.95}^{\pm{0.02}}$
[…]
@@ 11. Discussion and Future Work
In this work, we present StableMotion, a novel framework for training motion cleanup models directly on unpaired, corrupted motion datasets. StableMotion employs a diffusion-based generate-discriminate model that simultaneously evaluates motion quality and generates motions of varying quality through the use quality indicator variables. Our model can be trained with a simple masked inpainting strategy, and can then effectively clean a wide range of motion artifacts without relying on training with clean motion data, which is often costly and time-consuming to acquire. Experiments on SoccerMocap demonstrate StableMotion’s effectiveness in large-scale real-world production datasets.
Our framework demonstrates that a clean motion dataset is not necessarily a prerequisite for developing effective motion cleanup models, opening potential opportunities of training models on large-scale, raw motion data commonly found in animation and related fields. However, the inpainting procedure utilized in our framework can lead to drastic changes of the content of a motion clip in some scenarios. For applications with higher motion precision requirements, such as dexterous manipulation, test-time classifier guidance may be incorporated to better preserve the content of the original motions. Another notable limitation of our framework is its reliance on motion quality labels for training, which can still involve a considerable amount of manual effort to annotate. However, we show in Appendix D that effective motion cleanup models can be trained when only a small portion of the dataset is annotated with quality labels. Developing automated techniques that can accurately identify motion artifacts can further improve the effectiveness of our framework and expand its application across more diverse domains.
Acknowledgements.
This work was supported by NSERC (RGPIN-2015-04843) and the National Research Council Canada (AI4D-166). We would like to thank the anonymous reviewers for their constructive comments, Alessandro Sestini, Chuan Guo, and Shihao Zou for their support for this work.
@@ Appendix A Evaluation Metrics Details
  • Frozen Rate and Pops Rate: “Frozen frame” artifacts refer to segments where a character or specific joints remain unnaturally static for several consecutive frames, breaking the expected fluidity of motion. These often result from tracking failures, occlusions, or post-processing errors. “Motion pops” describe sudden, unnatural discontinuities or abrupt jumps in joint positions or orientations between frames, producing visually jarring transitions. We use a proprietary artifact detection algorithm, developed in a game studio production workflow, to compute Frozen Rate and Pops Rate, which quantify the frequency of frozen frames & over-smoothed transitions (Rempe et al., 2021), and motion pops, respectively.
@@
Figure 12. Our StableMotion model is implemented using a diffusion transformer (Peebles and Xie, 2023). The key modifications we introduce is the incorporation of rotary positional embeddings (RoPE) instead of absolute positional embeddings, which are applied to the hidden vectors $Q$ and $K$ in the multi-head self-attention module. The operation “@@” denotes application of a rotation to the vectors. The final key/query used for scaled dot-product attention is constructed by concatenating the original $Q/K$ with their rotated counterparts, $Q_{\text{rot}}/K_{\text{rot}}$.
Recall at Human-Annotated Artifacts (R@HAA). As described in Section 8, challenging artifacts in SoccerMocap, such as body penetration and subtle pose errors that are difficult for automatic algorithms to detect, are manually annotated by artists. We calculate our model’s artifact detection recall of these human-annotated artifacts (HAA). R@HAA is calculated following the recall definition:
(5) $\textbf{R@HAA}:=\frac{TP_{\text{HAA}}}{TP_{\text{HAA}}+FN_{\text{HAA}}},$